OpenStudy (anonymous):

(2v)/(v+1)=(-8)/(v^2-4v-5)

jhonyy9 (jhonyy9):

2v -8 ----- = ---------- ------ 2v(v-5)=-8 ------ 2v^2 -10v +8 =0 /2 v+1 (v-5)(v+1) v^2 -5v +4=0 (v-4)(v-1)=0 v_1 =4 v_2 =1